you are to choose between 2 different bands. flyer fever charges a rate of $470 and them $1.25 per ticket to play for the evening. eaglette charges $350 plus 2.00 for each ticket sold
write a linear equation showing the costs for each band. graph each of your lines to determine when the costs of the bands will be equal. use the elimination method to verify the solution from your graph is correct

Answer by nerdybill(7384)

you are to choose between 2 different bands. flyer fever charges a rate of $470 and them $1.25 per ticket to play for the evening. eaglette charges $350 plus 2.00 for each ticket sold
write a linear equation showing the costs for each band. graph each of your lines to determine when the costs of the bands will be equal. use the elimination method to verify the solution from your graph is correct
.
Let x = number of tickets sold
then
flyer fever:
cost = 1.25x + 470
eaglette:
cost = 2.00x + 350
.
to find when they're equal set the two equations equal to each other:
1.25x + 470 = 2.00x + 350
1.25x + 120 = 2.00x
120 = 0.75x
120/0.75 = x
160 tickets = x
(cost is equal if they sell 160 tickets)